If face value of equity shares of both companies ₹10 each, calculate the Earnings per share ratio for both companies, advising which company is recommended for investment. 


1
Expert's answer
2022-03-03T09:45:01-0500

Company A = 5/10 = 0.5

Company B = 6/10 = 0.6

Based on the above calculations, it is recommended to invest in the Company B.